40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Apr 26 01:00:23 2017, gautam, Update, General, DRMI locking  DRMI_1f_26April2017.pngIMG_7373.JPGIMG_7372.JPG
    Reply  Fri Apr 28 02:04:36 2017, gautam, Update, General, DRMI locking  20170428_DRMI.pdf
       Reply  Fri Apr 28 13:56:26 2017, rana, Update, General, DRMI locking  
          Reply  Mon May 1 16:29:51 2017, gautam, Update, General, DRMI locking  LSC_sensingNoise.pdfIMG_7378.JPG
             Reply  Mon May 1 21:45:54 2017, ericq, Update, General, DRMI locking  
                Reply  Wed May 3 18:45:45 2017, rana, Update, General, DRMI locking  
                   Reply  Thu May 4 19:03:15 2017, gautam, Update, General, DRMI locking - preliminary MICH NB DRMI_noArms_April30.pdfMICH_OLTF.pdfC1NB_disp_40m_MICH_NB_30_April_2017.pdf40m_REFL_RFPDs_efficiency.pdf
                      Reply  Fri May 5 10:13:02 2017, ericq, Update, General, MICH NB questions 
                         Reply  Fri May 5 12:10:53 2017, gautam, Update, General, MICH NB questions C1NB_disp_40m_MICH_NB_30_April_2017.pdf
                            Reply  Sat May 6 21:52:11 2017, rana, Update, General, MICH NB questions 
                               Reply  Wed May 10 01:56:06 2017, gautam, Update, General, MICH NB - OL coupling OL_BS_ITM_comp.pdfC1NB_disp_40m_MICH_NB_8_May_2017.pdfC1NB_disp_40m_MICH_NB_10_May_2017.pdf
                                  Reply  Wed May 10 16:53:38 2017, rana, Update, General, MICH NB - OL coupling 
                                  Reply  Wed May 10 17:17:05 2017, gautam, Update, General, DAC / Coil Driver noise actuation.jpg
                                     Reply  Wed May 10 17:46:44 2017, gautam, Update, General, DAC / Coil Driver noise - SRM coil driver + dewhite board removed 
                                        Reply  Thu May 11 09:45:46 2017, rana, Update, General, DAC / Coil Driver noise - SRM coil driver + dewhite board removed 
                                           Reply  Thu May 11 18:59:22 2017, gautam, Update, General, SRM coil driver + dewhite board initial survey D010001-B_40m.pdfD000183-C8_40m.pdf
                                              Reply  Fri May 12 01:36:04 2017, gautam, Update, General, SRM coil driver + dewhite board LISO modeling SRM_bypass_plus_CoilDriver.pdfliso.zip
                                                 Reply  Fri May 12 12:34:55 2017, gautam, Update, General, ITM and BS coil driver + dewhite board pulled out 
                                                    Reply  Fri May 12 18:50:08 2017, gautam, Update, General, ITM and BS coil driver + dewhite board pulled out 
                                                       Reply  Mon May 22 13:37:01 2017, gautam, Update, General, DAC noise estimate DAC_noise_model.pdfC1NB_disp_40m_MICH_NB_22_May_2017.pdf
                                                          Reply  Tue May 23 22:58:23 2017, gautam, Update, General, De-Whitening board noises ITMX_deWhite_ch1_noise.pdf
                                                             Reply  Thu May 25 19:27:29 2017, gautam, Update, General, Coil driver board noises  coilDriverNoises.pdf
                                                                Reply  Mon May 29 16:47:38 2017, gautam, Update, General, Coil driver boards reinstalled deWhitening_consolidated.pdfdeWhitening_measurements.zipETMY_OL.png
                                                                   Reply  Tue May 30 16:02:59 2017, gautam, Update, General, Coil driver boards reinstalled OL_noiseInjection.pdf
                                                                      Reply  Thu Jun 1 00:10:15 2017, gautam, Update, General, Coil driver boards reinstalled 
       Reply  Fri Apr 28 19:32:06 2017, gautam, Update, General, DRMI locking - PRCL angular FF PRC_angularFF.pdfPRC_TFs.pdf
Message ID: 12972     Entry time: Thu May 4 19:03:15 2017     In reply to: 12969     Reply to this: 12974
Author: gautam 
Type: Update 
Category: General 
Subject: DRMI locking - preliminary MICH NB 

Summary:

I've been playing around with Evan's NB code trying to put together a noise budget for the data collected during the DRMI locks last week. Here is what I have so far.

Attachment #1: Sensing matrix measurement.

  • This is basically to show that the MICH error signal is mostly in AS55Q.
  • The whitening gain used was 0dB, and the demod phase was -82 degrees.
  • The MICH sensing response was 5.31*10^8 V/m, where V is the demod board output. The 40m wiki RFPD page for AS55 says the RF transimpedance is ~550ohms, and I measured the Demod Board puts out 5.1V of IF signal (measured at after the Preamp, which is what goes to the ADC) for 1V of RF signal at the PD input. Using these numbers, and assuming a PD responsivity of 0.8 A/W at 1064nm, the sensing response is 2.37*10^5 W/m. I don't have a feeling yet for whether this is a reasonable number, but it would be a number to compare to what my Finesse model tells me to expect, for example.
  • Actuator calibration used to arrive at these numbers was taken from this elog

Attachment #2: MICH OLTF measurement vs model

  • In order to build the MICH OLTF model, I used MATLAB to put together the following transfer functions:
    • BS pendulum
    • Digital servo filters from LSC_MICH
    • Violin mode filters 
    • Analog/Digital AA and AI filters. For the digital AA/AI filters, I took the coefficients from /opt/rtcds/rtscore/release/src/fe/controller.c
  • The loop measurement was taken with digital filter modules FM1, FM2, FM3, FM7, FM9 engaged. 
  • In order to fit the model to the measurement, I tried finding the best-fit values for an overall loop gain and delay. 
  • The agreement between model and measurement isn't stellar, but I decided to push ahead for a first attempt. This loop TF was used to convert various noises into displacement noise for plotting.

Attachment #3: Noise budget

  • It took me a while to get Evan's code going, the main changes I made were to use nds2 to grab data instead of GWPy, and also to replace reading in .txt files with importing .mat files. This is a work in progress.
  • Noises plotted:
    • Measured - I took the in loop error signal and estimated the free-running displacement noise with the model OLTF, and calibrated it into metres using the sensing response measurement. This looks consistent with what was measured back in Dec 2015.
    • Shot noise - I used the measured DC power incident on the PD, 13mW, RF transimpedance of 550 V/A, and the V/m calibration factor mentioned above, to calculate this (labelled "Quantum Noise").
    • Dark noise - measured with PSL shutter closed.
    • Seismic noise, thermal noise, gas noise - calculated with GWINC

I think I did the various conversions/calibrations/loop algebra correctly, but I may have overlooked something. Now that the framework for doing this is somewhat set up, I will try and put together analogous NBs for PRCL and SRCL. 

GV 22 August 2017: Attachment #4 is the summary of my demod board efficiency investigations, useful for converting sensing measurement numbers from cts/m to W/m.

Attachment 1: DRMI_noArms_April30.pdf  69 kB  | Hide | Hide all
DRMI_noArms_April30.pdf
Attachment 2: MICH_OLTF.pdf  160 kB  | Hide | Hide all
MICH_OLTF.pdf
Attachment 3: C1NB_disp_40m_MICH_NB_30_April_2017.pdf  770 kB  | Hide | Hide all
C1NB_disp_40m_MICH_NB_30_April_2017.pdf
Attachment 4: 40m_REFL_RFPDs_efficiency.pdf  27 kB  Uploaded Tue Aug 22 15:36:30 2017  | Hide | Hide all
40m_REFL_RFPDs_efficiency.pdf
ELOG V3.1.3-